# Wasserstein–Aitchison GAN for angular measures of multivariate extremes

### v1

@@ Appendix C Angular measure with respect to another norm
[…]
@@
[…]
@@
$d\int_{\Delta_{d-1}}\int_{0}^{\infty}f(r\bm{w})\frac{\mathrm{d}r}{r^{2}}\,\mathrm{d}\Phi(\bm{w})=\int_{\widetilde{\Delta}_{d-1}}\int_{0}^{\infty}f(r\widetilde{\bm{w}})\frac{\mathrm{d}r}{r^{2}}\,\mathrm{d}\Psi(\widetilde{\bm{w}}),$
@@
$d\int_{\Delta_{d-1}}g\left(\frac{\bm{w}}{\|\bm{w}\|}\right)\|\bm{w}\|\mathrm{d}\Phi(\bm{w})=\int_{\widetilde{\Delta}_{d-1}}g(\widetilde{\bm{w}})\mathrm{d}\Psi(\widetilde{\bm{w}}).$
@@
$\int_{\widetilde{\Delta}_{d-1}}g(\widetilde{\bm{w}})\,\mathrm{d}\widetilde{\Phi}(\widetilde{\bm{w}})$ $\approx\frac{\int_{\Delta_{d-1}}g\left(\frac{\bm{w}}{\|\bm{w}\|}\right)\|\bm{w}\|\,\mathrm{d}\widehat{\Phi}(\bm{w})}{\int_{\Delta_{d-1}}\|\bm{w}\|\,\mathrm{d}\widehat{\Phi}(\bm{w})}$
$=\sum_{i=1}^{n}\Lambda_{i}g\left(\frac{\bm{\Theta}_{i}}{\|\bm{\Theta}_{i}\|}\right)=\int_{\widetilde{\Delta}_{d-1}}g(\widetilde{\bm{w}})\,\mathrm{d}\widehat{\widetilde{\Phi}}(\widetilde{\bm{w}}),$
@@ Appendix D Implementation of competing methods
HTGAN is follows a standard GAN architecture [19] with heavy-tailed latent variables that are Pareto($\alpha$) distributed for some $\alpha>0$. It does not make use of (multivariate) extreme value theory at all and rests on the capacity of the GAN to accurately capture the dependence of the data, also in the tail. As advised in their paper, we base our implementation on the one in [29]. Their preprocessing step in their Algorithm 1 on page 10, equation (3.11), requires an estimate of $\gamma=1/\alpha$ which is not provided in their methodology. They use the true value in their simulations, which is why we will also do so here, in Section 4.2. For real data applications, we will consider $\gamma=1/1.5$ as recommended by the authors in their results on page 19. Note also that it is not clear to us how step (3.12) in their Algorithm 2 is supposed to work as $\tilde{X}_{j}^{(i)}$ is not an element of $[0,1]$ so that it does not make sense to apply the inverse of the estimated marginal distribution function $F_{j}$. Since in this simulation setting the output of their generator is supposed to live in the same space as the original data, we do not apply any transform to compute the score (23). The score (22) is, of course, not affected by any marginal transformation. In real data applications, we do not compute the $W_{2}$ score and only consider the dependence score as done in the original paper. In terms of architectures and hyperparameters, we consider the same setting as for our method. This is very close to what the authors also do in their paper, besides the hyperparameter search which is performed through a Bayesian search in their numerical section. Both the generator and the discriminator are trained on $5000$ epochs. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures.
The GPGAN is closer to our approach as it is also based on the MGP distribution to accurately sample from the tail of $\bm{X}$, even though the approach in [28] is quite different. The authors also rely on a GAN architecture so that we also rely on [29] to implement their methodology. The main differences between their approach and ours are as follows:
  • In [28], the dependence is captured by the spectral vector $\bm{S}$ as introduced in (11), while our techniques rely on sampling from $\Phi$ after a transformation to the Aitchison coordinates which helps simplifying the distributions.
@@
The adaptive procedure in [28] to pick the threshold $\bm{u}$ based on the “extremeness” function $E$ was not entirely clear to us. Therefore, we decided to pick the same $\bm{u}$ as we do to compute the marginal parameters, so that we work with the same estimates of the parameters $\bm{\sigma}$ and $\bm{\xi}$. Also, as in [28] the latent distribution is not indicated, we consider a multivariate standard normal distribution on the latent space. The architectures and hyperparameters are also treated in the same way as for our method and the HTGAN method. Both the generator and the discriminator are trained on $5000$ epochs, based on the Adam optimizer, as done in the original paper [28]. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures. Note that computing the score (23) poses no problem with this method as it outputs directly on the scale of $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$; however, for the score (22), one needs to be more cautious. For this, we rely on the fact that if $\bm{X}^{*}_{1},\ldots,\bm{X}^{*}_{n}$ forms a sample from $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$, then
@@
The estimator can be plugged in the formula of extremal coefficient (21) which is used in computing the score (22).
@@ Appendix E Additional figures
Figure 9: GPD-qqplots for margins 1 to 10 in the financial dataset.
Figure 10: GPD-qqplots for margins 11 to 20 in the financial dataset.
Figure 11: GPD-qqplots for margins 21 to 30 in the financial dataset.

### v2

@@ Appendix C Angular measure with respect to another norm
The angular measure $\Phi$ in (2) can be introduced with respect to an arbitrary norm $\|\cdot\|$ on $\mathbb{R}^{d}$, that is, if we let
@@
for any Borel set $A\subseteq\widetilde{\Delta}_{d-1}$ such that $\widetilde{\Phi}(\partial A)=0$. Popular choices for $\|\cdot\|$ consist of the $L_{p}$ norms $|\cdot|_{p}$ on $\mathbb{R}^{d}$ for $p\in[1,\infty]$, see, e.g., [einmahl2001nonparametric, einmahl2009maximum, clemencon2023concentration]. Even though we only consider the norm $|\cdot|_{1}$ in Algorithm 1, we show that a simple post-processing step permits to provide an estimate for $\widetilde{\Phi}$ also.
@@
[…]
@@
[…]
@@
[…]
@@ Appendix D Implementation of competing methods
HTGAN is follows a standard GAN architecture [goodfellow2014] with heavy-tailed latent variables that are Pareto($\alpha$) distributed for some $\alpha>0$. It does not make use of (multivariate) extreme value theory at all and rests on the capacity of the GAN to accurately capture the dependence of the data, also in the tail. As advised in their paper, we base our implementation on the one in [github_GAN]. Their preprocessing step in their Algorithm 1 on page 10, equation (3.11), requires an estimate of $\gamma=1/\alpha$ which is not provided in their methodology. They use the true value in their simulations, which is why we will also do so here, in Section 4.2. For real data applications, we will consider $\gamma=1/1.5$ as recommended by the authors in their results on page 19. Note also that it is not clear to us how step (3.12) in their Algorithm 2 is supposed to work as $\tilde{X}_{j}^{(i)}$ is not an element of $[0,1]$ so that it does not make sense to apply the inverse of the estimated marginal distribution function $F_{j}$. Since in this simulation setting the output of their generator is supposed to live in the same space as the original data, we do not apply any transform to compute the score (23). The score (22) is, of course, not affected by any marginal transformation. In real data applications, we do not compute the $W_{2}$ score and only consider the dependence score as done in the original paper. In terms of architectures and hyperparameters, we consider the same setting as for our method. This is very close to what the authors also do in their paper, besides the hyperparameter search which is performed through a Bayesian search in their numerical section. Both the generator and the discriminator are trained on $5000$ epochs. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures.
The GPGAN is closer to our approach as it is also based on the MGP distribution to accurately sample from the tail of $\bm{X}$, even though the approach in [Li-et-al] is quite different. The authors also rely on a GAN architecture so that we also rely on [github_GAN] to implement their methodology. The main differences between their approach and ours are as follows:
  • In [Li-et-al], the dependence is captured by the spectral vector $\bm{S}$ as introduced in (11), while our techniques rely on sampling from $\Phi$ after a transformation to the Aitchison coordinates which helps simplifying the distributions.
@@
The adaptive procedure in [Li-et-al] to pick the threshold $\bm{u}$ based on the “extremeness” function $E$ was not entirely clear to us. Therefore, we decided to pick the same $\bm{u}$ as we do to compute the marginal parameters, so that we work with the same estimates of the parameters $\bm{\sigma}$ and $\bm{\xi}$. Also, as in [Li-et-al] the latent distribution is not indicated, we consider a multivariate standard normal distribution on the latent space. The architectures and hyperparameters are also treated in the same way as for our method and the HTGAN method. Both the generator and the discriminator are trained on $5000$ epochs, based on the Adam optimizer, as done in the original paper [Li-et-al]. The final layer of the discriminator has a sigmoid activation function, as it is commonly applied in GAN architectures. Note that computing the score (23) poses no problem with this method as it outputs directly on the scale of $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$; however, for the score (22), one needs to be more cautious. For this, we rely on the fact that if $\bm{X}^{*}_{1},\ldots,\bm{X}^{*}_{n}$ forms a sample from $\bm{X}\mid\bm{X}\nleq\bm{u}(t)$, then
@@ Appendix E Additional figures
(a) $d=10$
(b) $d=20$
(c) $d=50$
Figure 11: Heatmaps of the variogram matrices used in the simulations of Section 4.2.2.
Figure 12: Boxplots of estimated marginal GP parameters
Figure 13: GP qq-plots for margins 1 to 10 in the financial dataset.
Figure 14: GP qq-plots for margins $1$ to 20 in the financial dataset.
Figure 15: GP qq-plots for margins 21 to 30 in the financial dataset.